I'm having problems doing this word problem. Somebody help! Here it is:
A pharmacist mixed some 10% saline solution with some 15% saline solution to obtain 100ml of a 12% saline solution. How much of the 10% saline solution did the pharmacist use in the mixture?
Now, here is my work that i did. I made a 3 variable equation and something seems to not work there. Could you please show me what i did wrong or come up with a new way to approach the problem?
x+y=z
10x+15y= 100(12z)
Thanks a ton!

Answer by ikleyn(52786)

.
I'm having problems doing this word problem. Somebody help! Here it is:
A pharmacist mixed some 10% saline solution
with some 15% saline solution to obtain 100ml of a 12% saline solution. How much of the 10% saline solution did the pharmacist use in the mixture?
Now, here is my work that i did. I made a 3 variable equation and something seems to not work there. Could you please show me what i did wrong or come up with a new way to approach the problem?
x+y=z
10x+15y= 100(12z)
Thanks a ton!
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Your equations are (or should be)

  x +   y = 100,      (1)
10x + 15y = 12*100.   (2)


NO THREE EQUATIONS !   NO THREE UNKNOWNS !   Only two. 


Could you complete the solution from this point?


There are many similar solved problems in this site that you can use as samples.
See the lesson  Solving typical word problems on mixtures for solutions  in this site.